IN Resonance circuit If L=6H, C= 7 mF,

angular resonance frequency (ws) is:


Select one:

a. ws= 5.8 rad/sec

b. ws= 4.88 rad/sec

C. ws= 6.8 rad/sec NIC

d. ws = 8.9 rad/sec


1
Expert's answer
2021-06-07T05:49:40-0400

Angular frequency is given by formulae,

"\omega=\frac{1}{\sqrt{LC}}"

Substituting given values,

"\omega=\frac{1}{\sqrt{6\times7\times10^{-3}}}=4.88 rad/sec"


Hence option B is the correct answer.